Make the variable in the bracket the subject of the given formula
1. a= (1+b)/(1-b) {b}
1
Expert's answer
2012-09-07T11:10:23-0400
a= (1+b)/(1-b) {b}
a(1-b)=1+b
a-ab=1+b
a-1=b+ab
b(a+1)=a-1
b=(a-1)/(a+1)